OpenStudy (anonymous):

5th square root how do you solve them?

OpenStudy (anonymous):

You mean the 5th root? You generally plug it into your calculator ;p

OpenStudy (anonymous):

You can do it by approximating, but it's long and tedious.

OpenStudy (anonymous):

the following equation \[\sqrt[5]{64t^{17}u^{15}}\]

OpenStudy (anonymous):

sorry. what i mean is how to simplify them.

OpenStudy (anonymous):

Oh. Just recall that: \[\sqrt[a]{b^k} = b^{k\over a}\] And simplify the fractions in your exponents.

OpenStudy (anonymous):

So for example: \[64 = 2^6 \implies \sqrt[5]{64} = \sqrt[5]{2^6} = \sqrt[5]{2^5\cdot 2} = 2\sqrt[5]{2}\]

OpenStudy (anonymous):

polpak i get that. i don't get how to do the exponents.

OpenStudy (anonymous):

I just explained.. \[t^{17} = t^{15}\cdot t^2 \implies \sqrt[5]{t^{17} } = \sqrt[5]{t^{15} \cdot t^2} = t^{15/5}\sqrt[5]{t^2} = t^3\sqrt[5]{t^2}\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yeah that makes sense.

OpenStudy (anonymous):

ok. so this is what i got for the answer. \[2t^3u \sqrt[5]{2t^2}\] when simplified the radical form.

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Wait, no

OpenStudy (anonymous):

appreciate it polpak.

OpenStudy (anonymous):

Should be \(u^3\)

OpenStudy (anonymous):

sorry my bad. I wrote the problem wrong. since it should be \[\sqrt[5]{64t^{17}u^5}\] ends up to be correct. small mistake on my end.

OpenStudy (anonymous):

Oh I see. Yes.